A forward Horizontal force of 12N is used to pull a 240 N crate at a constant velocity across a horizontal floor. The coefficient of friction is?


1
Expert's answer
2021-03-23T11:11:36-0400

"F=ma"


"a=0,\ v=const"


"F-F_f=0\to F_f=F\to \mu N=F\to\mu=F/N"


"\mu=\frac{F}{N}=\frac{F}{mg}=\frac{12}{240}=0.05" . Answer